Enter your answer in the provided box. Calculate the pH of a 0.25 M methylamine solution. pH =
Answer #1

= CH₃NH₂ + H2O 0.25M (initial) CM₂ NH3t ton 0 0 0.250 0.25% 0.2501-4) Kb = (0,250) (0.250) = 0.25x² 0.25(1-x) 1-da 1 (since &x²= 4.4x154 = 17.6x104 0.25 d= (17.6 x 10-4) ? 1 : 4.20 X 100 [or ] = 0.258 = 0.25 X4.20x10*M = 1.05X16X Since; Po H = -log